OpenStudy (anonymous):

tan^2 theta-49=0....two solutions possible?

OpenStudy (anonymous):

Assuming you're solving for theta: \[\tan ^2(\theta-49)=\frac{ \sin (\theta-49) }{ \cos(\theta-49) }\] This = 0 when \[\sin(\theta-49)=0\] \[\sin(x)=0\] when \[x=0,\pi,2\pi,etc\] so \[\theta=0, \pi+49, 2\pi+49, etc\]

OpenStudy (noelgreco):

(tan x -7)(tan x +7)= 0